OpenStudy (anonymous):

Graph the function below. Determine the domain, range, and horizontal asymptote. Please show all of your work. f(x)=(3/4)^(x+3)

OpenStudy (anonymous):

Domain: \[\left\{ x| x \in \mathbb{R}\right\}\] Range: \[\left\{ y| y \in (0, \infty) \right\}\] Horizontal asymptote is: \[y=0\]

OpenStudy (anonymous):